OpenStudy (anonymous):

11/2 as a mixed number?

OpenStudy (anonymous):

What is 11 divided by 2? That is, how many times does 2 go into 11?

OpenStudy (anonymous):

1 1/2

OpenStudy (anonymous):

\[1\frac{1}{2}\] ? That's already a mixed number.

OpenStudy (anonymous):

that is what i needed. 11/2 as a mixed number is 1 1/2 correct?

OpenStudy (anonymous):

Yes. Mixed number- \(\Large2\frac{2}{3}\) improper fraction- \(\Large \frac{8}{3}\)

OpenStudy (anonymous):

@candire, just to be clear... \[\huge \frac{11}{2}\ne 1\frac{1}2{} \]

OpenStudy (anonymous):

* @candrie , ...

OpenStudy (anonymous):

Thank you. I had a blonde moment. I made it harder than it is